Find real numbers a and b such that:
(a+2i)(1+bi)=17-19i
(i as in imaginary)
Thank you :D

Answer by Jk22(389)

(a+2i)(1+bi)=a-2b+i(2+ab)=17-19i

a-2b=17, 2+ab=-19, ab=-21, 21=3*7, try

a b---: a-2b
-3 7--:-3-14=-17

the solution is a=-3, b=7